Suppose that one-way commute times in a particular city are normally distributed with a mean of 19.32 minutes and a standard deviation of 2.364 minutes. Would it be unusual for a commute time to be above 22.9 minutes?

Question 6 options: 1) The value is unusual. 2) It is impossible for this value to occur with this distribution of data. 3) We do not have enough information to determine if the value is unusual. 4) The value is not unusual. 5) The value is borderline unusual.
Answer #1

M = 19:32, o = 2364 P(x9 2209 ) = P (Z) 22-9-19032 2.36 4 0.0655 = P(2) 1.51) -- P(Z <1.51). = - 0.9345 = [0-06557 Alios * T